OpenStudy (anonymous):

Can someone please explain how to do problems such as... (4 - 3sqrt7)[2 + 5sqrt7] ?

OpenStudy (andrewkaiser333):

what is the lesson name and chapter then i can give you some ways of helping

OpenStudy (andrewkaiser333):

learn it well

OpenStudy (anonymous):

I'm actually doing math placement review so I don't have an exact name, sorry.

OpenStudy (anonymous):

thank you!

OpenStudy (andrewkaiser333):

tat will help you learn it, it is similar to yours

OpenStudy (andrewkaiser333):

that* ok np that shall help

OpenStudy (andrewkaiser333):

tell me if you still do not understand

OpenStudy (mathstudent55):

You are multiplying two numbers. Each number has a rational part and an irrational part. Multiply the numbers using FOIL, then combine like terms.

OpenStudy (andrewkaiser333):

here this helps too they just went through one similar using foil. http://openstudy.com/study#/updates/5175f782e4b0be9997e05dd0

OpenStudy (anonymous):

this is probably incorrect but is the answer to the example above -97 + 14 sqrt7?

OpenStudy (mathstudent55):

Let me check.

OpenStudy (mathstudent55):

\((4 - 3 \sqrt 7) (2 + 5 \sqrt 7) \) \(= 8 + 20 \sqrt 7 - 6 \sqrt 7 - 15 (\sqrt 7)^2 \) \(= 8 + 14 \sqrt 7 - 15 \cdot 7 \) \(= 8 + 14 \sqrt 7 - 105 \) \(= -97 + 14 \sqrt{7} \) This is correct as a final answer \(=14\sqrt 7 - 97\) and so is this Correct. Good job!

OpenStudy (anonymous):

thank you so much!!!

OpenStudy (mathstudent55):

You are very welcome.